A business started 1 April 2014, and incurred the following costs during its first two years.

Year ending 31 March  2014    2015

           

Direct materials            60,000 49,900

Direct labour               48,000 44,000

Variable overheads       24,000 30,000

Fixed costs                 40,000 40,600             

 

Production each year  16,000    14,000

Sales each year    14,000          14,000

Do:

Prepare a statement showing the gross profit for each of the three years if the company used:

The marginal costing approach to valuing  inventory;   

The absorption costing approach to valuing inventory. 

Advantages and disadvantages of using each method


1
Expert's answer
2021-12-05T18:55:26-0500

The marginal costing:

"(60,000+48,0000+24,000)\/16,000*14000=115,500" 2014 year

"8,25*2,000+(49,000+44,000+30,000)\/14,000*12,000=121,928.57" 2015 year

This method allows you to estimate the costs associated directly with the production of the product, which allows you to understand how much the product itself costs. But this method does not allow you to estimate the indirect costs in the value of the goods.

The absorption costing

"(60,000+48,0000+24,000+40,000)\/16,000*14000=150,500" 2014 year

"10,75*2,000+(49,000+44,000+30,000+40,600)\/14,000*12,000=161,728.57" 2015 year

This method allows you to carry forward all costs for a product, but it underestimates the gross margin, which makes it impossible to understand the gross margin for a particular product. If there is more than one type of production, then this method can distort the gross profit of an individual product and lead to incorrect conclusions.